OpenStudy (anonymous):

Tommy's piggy bank contains $8.40 in nickels and pennies if there were 60 more nickels than pannies, how many of each did he have on the piggy bank?

OpenStudy (lgbasallote):

let n = nickels = 60 + p (so...p = n-60) p = pennies 5(cents)n + 1 (cent)p = 840 cents 5n + p = 840 so....p = 840 - 5n since it has been sated a while ago that p = n -60 and p = p then... n - 60 = 840 - 5n 5n + n = 840 + 60 6n = 900 n = 900/6 n = 150 p = 150 - 60 = 90

OpenStudy (anonymous):

Thank you very much!